OpenStudy (anonymous):

8- 6.2 divided by 5 + 7(0.91) help me please

OpenStudy (anonymous):

13.3

OpenStudy (anonymous):

My textbook says 13.13 so I was wondering how they got that. Can you show me/ Tell me how you got that answer because I'm kinda confused...

OpenStudy (anonymous):

its easy bc its basic order of operations

OpenStudy (anonymous):

Yeah I know NVM :) thanks

OpenStudy (anonymous):

can i hav a medal o.O

OpenStudy (anonymous):

tyvm

OpenStudy (anonymous):

np

OpenStudy (asnaseer):

for this:\[\frac{8- 6.2}{5 + 7(0.91)}\] you need to follow BODMAS - have you heard about that?

OpenStudy (asnaseer):

it is known by other names as well, see here: http://en.wikipedia.org/wiki/Order_of_operations

OpenStudy (anonymous):

Ok let me check it out hold on

OpenStudy (asnaseer):

look at the 2nd section "The standard order of operations"

OpenStudy (anonymous):

Ok hold on hmm SO MANY WORDS!! D:

OpenStudy (anonymous):

lol typical wikipedia

OpenStudy (anonymous):

Ok I kinda got it

OpenStudy (anonymous):

(not really so much)

OpenStudy (anonymous):

kwl i hope u pass :)

OpenStudy (asnaseer):

good - so first thing to attack is the brackets

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

^^...

OpenStudy (asnaseer):

in this case the brackets just contain a number - so no actual operation to perform inside the brackets

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

we can therefore remove the brackets to end up with:\[\frac{8- 6.2}{5 + 7(0.91)}=\frac{8- 6.2}{5 + 7*0.91}\]

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

Mhm ok

OpenStudy (anonymous):

hmmm ok im confused...

OpenStudy (asnaseer):

lets jus concentrate on the denominator for now, we had:\[5+7(0.91)\]correct?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

now note that a(b) simply means a*b - agreed?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

agreed

OpenStudy (asnaseer):

so we can say that:\[5+7(0.91)=5+7*0.91\]

OpenStudy (asnaseer):

does that make sense now?

OpenStudy (anonymous):

ok got tht

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good - now what should the next set of operations be?

OpenStudy (asnaseer):

we have handled the brackets

OpenStudy (anonymous):

um to multiply?>

OpenStudy (anonymous):

sorry didnt mean to put >

OpenStudy (asnaseer):

it should be exponents and roots next - but you do not have any here so we go to the next one: multiplication an division

OpenStudy (anonymous):

Yea!! Ok

OpenStudy (asnaseer):

so what is 7 * 0.91 = ?

OpenStudy (anonymous):

Um hold on

OpenStudy (anonymous):

6.37

OpenStudy (asnaseer):

correct, so now we end up with:\[\frac{8- 6.2}{5 + 7(0.91)}=\frac{8- 6.2}{5 + 7*0.91}=\frac{8- 6.2}{5 + 6.37}\]now most people get confused here because they think the next thing to do is the division

OpenStudy (anonymous):

wouldnt you add? or subtract?

OpenStudy (anonymous):

add?

OpenStudy (asnaseer):

but what you need to remember is that you go through BODMAS for the numerator and the denominator seperately

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

hang on a second - I think I misread your question

OpenStudy (asnaseer):

I believe it was this instead:\[8- 6.2 \div 5 + 7(0.91)\]

OpenStudy (asnaseer):

am I right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, sorry I originally interpreted it as:\((8- 6.2) \div (5 + 7(0.91))\)

OpenStudy (anonymous):

ohhh ok

OpenStudy (asnaseer):

so lets start again - ok?

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

\[8- 6.2 \div 5 + 7(0.91)\]first step - brackets to get:\[8- 6.2 \div 5 + 7(0.91)=8- 6.2 \div 5 + 7*0.91\]next is exponents/root but we have none next is multiplication/division - we have both, so you did 7*0.91 = 6.37\[8- 6.2 \div 5 + 6.37\]

OpenStudy (asnaseer):

thats how far we got

OpenStudy (asnaseer):

so next we should do what?

OpenStudy (anonymous):

add?

OpenStudy (anonymous):

or no subtract

OpenStudy (asnaseer):

nope - have a closer look at what is still left

OpenStudy (anonymous):

like left to right but switch add with division

OpenStudy (asnaseer):

it should be division next

OpenStudy (anonymous):

so divide?

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

so what is 6.2 divided by 5 = ?

OpenStudy (anonymous):

ok hold on

OpenStudy (asnaseer):

(a quick trick to divide by 5 is to divide by 10 and then halve the result)

OpenStudy (anonymous):